A ball is thrown vertically upwards with speed u. If it experiences a constant air resistance force of magnitude f, then the speed with which ball strikes the ground, (weight of ball is w) is

Grade:12

1 Answers

rishabh Tiwari
7 years ago
Let the ball have a mass `m` ,We have for upward motion :U=u ,V=0 ,A= -(g+f/m) = -( w+f )/m ,Using 3rd eq.n of motion we get:H = 0.5 *(mu^2)/( w+f )Now for downward motion we have:-A=(w-f)/m ,U =0 ,V= v(to find) ,Again use 3rd eq.n of motion to get : v = u*[sq.rt (w-f)/(w+f)]...ans.